OpenStudy (anonymous):

The perimeter of the sector is 124m. If sector angle of 36 degrees, then the radius is:

OpenStudy (anonymous):

I know the formula is \[2 \pi r + 2r\] can't seem to solve it though

OpenStudy (shamim):

i know a formula is theta=s/r

OpenStudy (shamim):

theta is in radian unit

OpenStudy (john_es):

If you mean, the length of the total sector, not only the curve part, then the formula cannot be 2 pi r+ 2r, because the angle is not 2 pi. You can correct the formula, \[\theta r+2r=124\] with theta in radians, \[\theta=36\pi/180\] And then solve for r.